OpenStudy (anonymous):

Okay I know this is really simple but for some reason I keep getting the wrong answer. I'm probably making some dumb math error in my math: Solve the following system of equations: 4x - 2y - z = -5 x - 3y + 2z = 3 3x + y - 2z = -5

OpenStudy (anonymous):

Oh these are the answer choices btw: A. (0, 1, 3) B. (0, -1, 3) C. (0, 1, -3) D. (0, -1, -3)

OpenStudy (anonymous):

Adding the 2nd and 3rd equations together seems to help.

OpenStudy (anonymous):

Because it leaves you with only two 4x-2y-z=-5 4x-2y=-2

OpenStudy (anonymous):

and since (4x-2y)=(-2), you can use that in the first equation (4x-2y)-z=-5 -2-z=-5

OpenStudy (anonymous):

z=3

OpenStudy (anonymous):

so plug that into -2-z=-5 and that gives you -5=-5?

OpenStudy (anonymous):

You can't plug an answer in an equation you just used to find that answer, you would just get a true statement. Plug it in the original equations.

OpenStudy (anonymous):

So I think the answer is B...

OpenStudy (anonymous):

x-3y+2*3=3 3x+y-2*3=-5 x-3y+3=0 3x+y-1=0 (y)=(1-3x) x-3(y)+3=0 x-3(1-3x)+3=0 10x=0 [x=0]

OpenStudy (anonymous):

Ahhhhh...got it!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No problemo :)